# Framing global structural identifiability in terms of parameter symmetries

### v1

@@
Framing global structural identifiability in terms of parameter symmetries
Johannes G Borgqvist [^1] Corresponding author: johborgq@chalmers.se [^2] Mathematical Sciences, Chalmers University of Technology, Gothenburg, Sweden, Alexander P Browning [^3] Mathematical Institute, University of Oxford, United Kingdom, Fredrik Ohlsson [^4] Department of Mathematics and Mathematical Statistics, Umeå University, Umeå, Sweden, Ruth E Baker3
A key initial step in mechanistic modelling of dynamical systems using first-order ordinary differential equations is to conduct a global structural identifiability analysis. This entails deducing which parameter combinations can be estimated from certain observed outputs. The standard differential algebra approach answers this question by re-writing the model as a system of ordinary differential equations solely depending on the observed outputs. Over the last decades, alternative approaches for analysing global structural identifiability based on so-called full symmetries, which are Lie symmetries acting on independent and dependent variables as well as parameters, have been proposed. However, the link between the standard differential algebra approach and that using full symmetries remains elusive. In this work, we establish this link by introducing the notion of parameter symmetries, which are a special type of full symmetry that alter parameters while preserving the observed outputs. Our main result states that a parameter combination is structurally identifiable if and only if it is a differential invariant of all parameter symmetries of a given model. We show that the standard differential algebra approach is consistent with the concept of considering structural identifiability in terms of parameter symmetries. We present an alternative symmetry-based approach, referred to as the CaLinInv-recipe, for analysing structural identifiability using parameter symmetries. Lastly, we demonstrate our approach on a glucose-insulin model and an epidemiological model of tuberculosis.
@@
Global structural identifiability, Parameter symmetries, Universal parameter invariants, Differential algebra approach, CaLinInv-recipe.
@@ 1 Introduction
Given an abundance of experimental data, a large focus of mechanistic modelling in biology concerns model validation and experimental design. A crucial initial step is to conduct a so-called structural identifiability analysis in order to deduce what model parameters can and cannot be estimated given experimental data. In the context of mechanistic models consisting of ordinary differential equations (ODEs) there is a plethora of methods (see [1] for a review). Importantly, the standard method for assessing so-called global structural identifiability of systems of ODEs depending on rational functions of the independent and dependent variables is referred to as the differential algebra approach [2, 3, 4, 5]. Essentially, this approach begins by re-formulating the original system of ODEs as an equivalent input-output system consisting of polynomial ODEs, of potentially much higher order, that depend solely on the observed inputs and outputs in addition to the rate parameters. After this re-formulation, the standard differential algebra approach entails constructing a map between the parameters in the original system of ODEs and the parameter combinations realised in the input-output system. Given such a parameter map, a global structural identifiability analysis entails assessing whether or not the parameter map is injective. If this is the case the model of interest is said to be structurally identifiable otherwise it is referred to as structurally unidentifiable. More specifically, the differential algebra-approach uses the fact that the input-output system is polynomial and it finds identifiable parameter quantities by extracting the coefficients in front of each monomial. In addition to the standard differential algebra approach, other methodologies for analysing the structural identifiability of systems of first-order ODEs have also been developed.
@@
For more than two decades, methods for analysing global structural identifiability using a special type of Lie point symmetries we refer to as full symmetries have been developed [6, 7, 8, 9, 10]. Full symmetries are transformations called $C^{\infty}$ diffeomorphisms which map solutions of the system of ODEs of interest to other solutions while simultaneously preserving the observed outputs. In particular, they act on the independent variable (corresponding to time), the dependent variables (corresponding to the states), the outputs and the parameters. An example of a type of full symmetry studied by Castro and de Boer [9] are scalings—where parameters and states are scaled by scaling factors that leave the system of interest invariant. Provided such scaling symmetries, parameters are identifiable and states are observable if they are invariant under scalings, which implies that their scaling factors all equal one [9]. Critically, not all models possess scalings as full symmetries and there are models with other types of full symmetries which can be used as a basis for deducing global structural identifiability as well [11]. Hence, solely relying on scaling symmetries in the context of global structural identifiability can be misleading [11].
Importantly, full symmetries constitute a generalisation of what we refer to as classical symmetries of ODEs which map solutions to other solutions by acting on the independent and dependent variables but not the outputs and the rate parameters. Notably, the main applications of classical symmetries are to find analytical solutions of ODEs, reduce the order of ODEs in order to present them in a simpler form, and to construct classes of ODEs from a set of symmetries [12, 13, 14, 15]. A well-known problem in the context of classical symmetries of first-order ODEs is that the dimension of the symmetry group of such systems is infinite. This implies that the so-called linearised symmetry conditions, which are the equations that must be solved in order to find the so-called infinitesimals (the functions that characterise symmetries), are always underdetermined in the case of system of first-order ODEs. In other words, there are more infinitesimals, i.e. unknowns, that characterise classical symmetries of first-order ODEs than there are linearised symmetry conditions, i.e. equations, and hence there is no straightforward methodology for finding the unknown infinitesimals except for constructing ansätze for them. This problem is even worse for full symmetries of first-order ODEs as the corresponding linearised symmetry conditions are yet more underdetermined as additional infinitesimals for parameters are introduced. Consequently, a large emphasis in global structural identifiability analyses based on full symmetries has been put on calculating full symmetries in an automated fashion [7, 8]. Technically, these approaches substitute ansätze for the unknown infinitesimals that are multivariate polynomials of the parameters, as well as the independent and dependent variables, into the linearised symmetry conditions which cause them to decompose into a system of linear equations that can be solved using Gaussian elimination. Nevertheless, there are still conceptual, fundamental and unanswered questions about the role of full symmetries in the context of global structural identifiability. Specifically, what is the exact link between global structural identifiability and full symmetries? Also, is the standard differential algebra approach consistent with the notion of full symmetries? If this is the case, this implies that we can use the powerful machinery of symmetry methods in order to deduce global structural identifiability of mechanistic models.
@@
In this work, we establish the link between global structural identifiability and full symmetries by introducing the notion of parameter symmetries. These are full symmetries solely acting on parameters, i.e. Lie symmetries acting as re-parametrisations of the model of interest. Provided the notion of parameter symmetries, we show that a parameter is globally structural identifiability if and only if it is a so-called universal parameter invariant which is a differential invariant of all parameter symmetries of a given model. Next, we demonstrate that the standard differential algebra approach for deducing global structural identifiability will always find universal parameter invariants and thus it is consistent with the notion of parameter symmetries. Thereafter, we develop an alternative methodology for deducing global structural identifiability based on parameter symmetries referred to as the CaLinInv-recipe (Algorithm 2). Specifically, the method: (i) re-writes the original system of ODEs in terms of the observed outputs referred to as the canonical coordinates (Ca); (ii) finds parameter symmetries by solving the linearised symmetry conditions (Lin); and (iii) calculates universal parameter invariants (Inv). Our approach finds both the globally structurally identifiable parameter quantities as well as the parameter transformations preserving these parameter quantities, namely the parameter symmetries. Lastly, we conduct global structural identifiability analyses of a glucose-insulin model and an epidemiological SEI model using the CaLinInv-recipe, yielding insights about identifiable parameter combinations as well as the family of parameter transformations which leaves them invariant. Overall, we demonstrate how global structural identifiability of mechanistic models can be understood in terms of parameter symmetries.
2 Mathematical preliminaries
@@
We briefly present the mathematical preliminaries of global structural identifiability in two parts, starting with the standard differential algebra approach and concluding with full symmetries. To this end, consider the following system of first-order ODEs and associated outputs
@@
$\begin{split}\dot{\mathbf{x}}&=\mathbf{f}\left(t,\mathbf{x},\boldsymbol{\theta}\right),\\ \mathbf{y}&=\mathbf{h}\left(t,\mathbf{x},\boldsymbol{\theta}\right),\end{split}$ (1)
@@
where derivatives are denoted by $\dot{\mathbf{x}}=\mathrm{d}\mathbf{x}/\mathrm{d}t$. Here, $t\in\mathbb{R}$ is the independent variable corresponding to time, $\boldsymbol{\theta}\in\mathbb{R}^{p}$ is the vector of $p$ parameters, $\mathbf{x}(t,\boldsymbol{\theta})\in\mathbb{R}^{n}$ are the $n$ dependent variables corresponding to the states of the system and $\mathbf{y}(t,\mathbf{x},\boldsymbol{\theta})\in\mathbb{R}^{m}$ are the $m\leq n$ observed outputs. In the context of global structural identifiability, we ask ourselves which of the $p\in\mathbb{Z}_{+}$ parameters collected in $\boldsymbol{\theta}\in\mathbb{R}^{p}$ can be identified based on a set of $1\leq m\leq n$ observed outputs $\mathbf{y}$ given $n\geq m$ states $\mathbf{x}$? Here, we make two important assumptions, namely that the functions $\mathbf{f}$ and $\mathbf{h}$ are analytical and rational functions of the independent and dependent variables and that they are infinitely differentiable. These assumptions are required to implement the standard differential algebra approach for analysing global structural identifiability whereas the symmetry-based approach can be carried out without them. However, since the the aim of this work is to establish the link between these two approaches we restrict ourselves to rational functions $\mathbf{f}$ and $\mathbf{h}$. Under these assumptions, we can always reduce the original system of first-order ODEs to an input-output system corresponding to a (potentially higher order) polynomial system of ODEs solely depending on the outputs [2]
@@
$\Delta\left(t,\frac{\mathrm{d}^{N}\mathbf{y}}{\mathrm{d}t^{N}},\frac{\mathrm{d}^{N-1}\mathbf{y}}{\mathrm{d}t^{N-1}},\ldots,\ddot{\mathbf{y}},\dot{\mathbf{y}},\mathbf{y},\boldsymbol{\theta}\right)=\mathbf{0},\$ (2)
for some power $N\in\mathbb{N}_{+}$ and where $\Delta$ is a vector-valued function of multivariate polynomials. Provided this problem formulation, we begin by presenting the definition of global structural identifiability as well as the standard differential algebra approach for conducting a global structural identifiability analysis. Thereafter, we present the notion of full symmetries in the context of structural identifiability.
2.1 Global structural identifiability and the standard differential algebra approach
We first present the definition of global structural identifiability [16].
@@ 2.1 Global structural identifiability and the standard differential algebra approach
An individual parameter $\theta_{j}\in\boldsymbol{\theta}$, ${j}\in{1,\ldots,p}$ is (globally) structurally identifiable if for almost every value $\boldsymbol{\theta}^{\star}$ and almost all initial conditions the following holds:

### v3

# Framing structural identifiability in terms of parameter symmetries

@@
Framing structural identifiability in terms of parameter symmetries
Johannes G Borgqvist [^1] Corresponding author: johborgq@chalmers.se [^2] Mathematical Sciences, Chalmers University of Technology, Gothenburg, Sweden, Alexander P Browning [^3] Mathematical Institute, University of Oxford, United Kingdom [^4] School of Mathematics and Statistics, University of Melbourne, Melbourne, Australia, Fredrik Ohlsson [^5] Department of Mathematics and Mathematical Statistics, Umeå University, Umeå, Sweden, Ruth E Baker3
A key step in mechanistic modelling of dynamical systems is to conduct a structural identifiability analysis. This entails deducing which parameter combinations can be estimated from a given set of observed outputs. The standard differential algebra approach answers this question by re-writing the model as a higher-order system of ordinary differential equations that depends solely on the observed outputs. Over the last decades, alternative approaches for analysing structural identifiability based on Lie symmetries acting on independent and dependent variables as well as parameters, have been proposed. However, the link between the standard differential algebra approach and that using full symmetries remains elusive. In this work, we establish this link by introducing the notion of parameter symmetries, which are a special type of full symmetry that alter parameters while preserving the observed outputs. Our main result states that a parameter combination is locally structurally identifiable if and only if it is a differential invariant of all parameter symmetries of a given model. We show that the standard differential algebra approach is consistent with the concept of structural identifiability in terms of parameter symmetries. We present an alternative symmetry-based approach for analysing structural identifiability using parameter symmetries. Lastly, we demonstrate our approach on two well-known models in mathematical biology.
@@
Local structural identifiability, parameter symmetries, universal parameter invariants, differential algebra approach.
@@ 1 Introduction
Given the abundance of experimental data, a substantial amount of effort in mechanistic modelling in biology goes into considering model validation and experimental design. A crucial preliminary step, prior to data collection, is to conduct a structural identifiability analysis to determine which model parameters can, in principle, be estimated from a given experimental setup. Structural identifiability is a theoretical property of a mechanistic model that can be investigated analytically under the assumption of perfect, noise-free observations of the model output. In this context, structural identifiability addresses whether parameters can be uniquely inferred from the observed output. A model parameter is said to be globally structurally identifiable if it can be uniquely determined, and locally structurally identifiable if it admits only finitely many distinct values consistent with the observed output.
For models formulated as systems of ordinary differential equations (ODEs), a wide range of methods for assessing structural identifiability is available (see [1] for a review). The standard approach for establishing global structural identifiability of ODE models with rational right-hand sides is based on differential algebra [2, 3, 4, 5]. This approach reformulates the original system as an equivalent input–output representation comprising polynomial differential equations, potentially of higher order, involving only the observed inputs, outputs, and model parameters. Structural identifiability is then assessed by constructing a map between the original model parameters and the parameter combinations appearing in the input–output equations, and determining whether this map is injective. Identifiable parameter quantities can subsequently be recovered by exploiting the polynomial structure of the input–output system, typically through analysis of the corresponding coefficients.
To illustrate the differential algebra approach, let $u(t)$ and $v(t)$ denote the concentrations of two chemical species that depend on time, $t$, and which satisfy two decoupled ODEs
$\dfrac{\mathrm{d}u}{\mathrm{d}t}=\kappa_{1}-\lambda{u},\quad\dfrac{\mathrm{d}v}{\mathrm{d}t}=\kappa_{2}-\lambda{v},$ (1)
where both species decay at rate $\lambda>0$ and are synthesised at rates rates $\kappa_{1}>0$ and $\kappa_{2}>0$, respectively. We assume that we observe only the total concentration $y(t)=u(t)+v(t)$, yielding the following model for the output $y$:
$\dfrac{\mathrm{d}y}{\mathrm{d}t}+\lambda{y}-(\kappa_{1}+\kappa_{2})=0.$ (2)
Next, we extract the coefficients in front of $\{\mathrm{d}y/\mathrm{d}t,y,1\}$ resulting in the set $\{1,\lambda,\kappa_{1}+\kappa_{2}\}$. Clearly, the set of identifiable parameter quantities is given by $\{\lambda,\kappa_{1}+\kappa_{2}\}$ implying that the decay rate $\lambda$ is (globally) structurally identifiable. On the other hand, the synthesis rates $\kappa_{1}$ and $\kappa_{2}$ are individually non-identifiable whereas the sum $\kappa_{1}+\kappa_{2}$ is globally structurally identifiable.
@@
In addition to the standard differential algebra approach, other methodologies for analysing the structural identifiability of systems of first-order ODEs have also been developed. In particular, for more than two decades, methods for analysing structural identifiability using a special type of Lie point symmetries we refer to as full symmetries have been developed [6, 7, 8, 9, 10]. Full symmetries are transformations called $C^{\infty}$ diffeomorphisms which map solutions of the system of ODEs of interest to other solutions while simultaneously preserving the observed outputs. In particular, they act on the independent variable (corresponding to time in our examples), the dependent variables (corresponding to the states), the outputs and the parameters.
An example of a class of full symmetries studied by Castro and de Boer [9] is scalings, in which state variables and parameters are transformed multiplicatively by scaling factors that leave the model equations invariant. Within this framework, Castro and de Boer analysed both the structural identifiability of parameters and the related notion of structural observability of states, defined as the ability to reconstruct state variables from observed outputs. In the presence of such scaling symmetries, parameters are structurally identifiable and states are structurally observable if and only if they are fixed points of the induced scaling action; equivalently, the only scaling transformation that preserves the observed outputs is the identity transformation, for which all scaling factors are equal to one. Critically, not all models possess scalings as full symmetries, and there are models with other types of full symmetries that can also be used as a basis for deducing structural identifiability [11]. Hence, relying solely on scaling symmetries in the context of structural identifiability can be misleading [11].
Importantly, full symmetries generalise the classical notion of symmetries of ordinary differential equations, which act on the independent and dependent variables and map solutions to solutions, but do not act on model parameters or observed outputs. Classical symmetry methods are primarily used to construct analytical solutions, reduce the order of ODEs, or classify families of differential equations with prescribed symmetry properties [12, 13, 14, 15]. A well-known limitation of classical symmetry analysis for first-order ODE systems is that their symmetry groups are infinite-dimensional. As a consequence, the associated linearised symmetry conditions—whose solutions define the infinitesimal generators of the symmetry group—are underdetermined, with more unknown infinitesimals than equations. In practice, this necessitates the introduction of restrictive ansätze in order to obtain tractable solutions. This difficulty is exacerbated in the context of structural identifiability analysis, where full symmetries introduce additional infinitesimals associated with model parameters, further increasing the degree of underdetermination. As a result, recent work on symmetry-based identifiability analysis has focused on the automated computation of full symmetries [7, 8]. These approaches impose polynomial ansätze on the infinitesimals—taken to be multivariate polynomials in the independent variables, state variables, and parameters—thereby reducing the linearised symmetry conditions to a finite-dimensional linear system that can be solved algorithmically, for example using Gaussian elimination. Despite these advances, fundamental questions remain unresolved, notably the precise relationship between full symmetries and structural identifiability, and the extent to which symmetry-based approaches are consistent with the classical differential algebra framework. Resolving these questions is essential if the full machinery of symmetry methods is to be systematically exploited for identifiability analysis in mechanistic models.
@@
In this work, we establish the link between structural identifiability and full symmetries by introducing the notion of parameter symmetries. These are full symmetries solely acting on parameters; that is, Lie symmetries acting as re-parametrisations of the model of interest. Provided the notion of parameter symmetries, we show that a parameter is locally structurally identifiable if and only if it is a so-called universal parameter invariant—a differential invariant of all parameter symmetries of a given model. Next, we demonstrate that the standard differential algebra approach for deducing global structural identifiability will always find universal parameter invariants and thus it is consistent with the notion of parameter symmetries. Subsequently, we develop an alternative methodology for deducing local structural identifiability based on parameter symmetries (Algorithm 1). Specifically, the method: (i) re-writes the original system of ODEs in terms of the observed outputs which are called the canonical coordinates; (ii) finds parameter symmetries by solving the linearised symmetry conditions; and (iii) calculates universal parameter invariants. Our approach finds both the locally structurally identifiable parameter quantities as well as the parameter transformations preserving these parameter quantities, namely the parameter symmetries. Lastly, we conduct local structural identifiability analyses of a glucose-insulin model and an epidemiological SEI model using Algorithm 1, yielding insights about identifiable parameter combinations as well as the family of parameter transformations which leaves them invariant. Overall, we demonstrate how local structural identifiability of mechanistic models can be understood in terms of parameter symmetries.
For the interested reader, we present definitions of global and local structural identifiability as well as the standard differential algebra approach in Section 2. Thereafter, we present the theoretical link between our parameter symmetries and local structural identifiability in Section 3. For the reader who is interested in the application of Lie symmetries in local structural identifiability analyses of models consisting of systems of ODEs, proceeding to application of parameter symmetries in local structural identifiability analyses of specific models in biology in Section 4 is sufficient.
2 Definitions of structural identifiability
@@
We briefly present definitions of both global and local structural identifiability followed by an algorithm for the standard differential algebra approach. To this end, consider the following state-output system of first-order ODEs and associated outputs
@@
$\begin{split}\dfrac{\mathrm{d}\mathbf{x}}{\mathrm{d}t}&=\mathbf{f}\left(t,\mathbf{x},\boldsymbol{\theta}\right),\\ \mathbf{y}&=\mathbf{h}\left(t,\mathbf{x},\boldsymbol{\theta}\right).\end{split}$ (3)
@@
Here, $t\in\mathbb{R}$ is the independent variable corresponding to time, $\boldsymbol{\theta}\in\mathbb{R}^{p}$ is the vector of $p$ parameters, $\mathbf{x}(t,\boldsymbol{\theta})\in\mathbb{R}^{n}$ are the $n\geq 1$ dependent variables corresponding to the states of the system and $\mathbf{y}(t,\mathbf{x},\boldsymbol{\theta})\in\mathbb{R}^{m}$ are the $1\leq{m}\leq n$ observed outputs. In the context of structural identifiability, we ask ourselves which parameters can be determined based on a set of $1\leq m\leq n$ observed outputs, $\mathbf{y}$, given $n$ states, $\mathbf{x}$? Here, we make two important assumptions, namely that the functions $\mathbf{f}$ and $\mathbf{h}$ are analytical and rational functions of the independent and dependent variables and that they are infinitely differentiable. Under these assumptions, we can always reduce the original state-output system of first-order ODEs to an output system corresponding to a (potentially higher order) polynomial system of ODEs solely depending on the outputs [2]
@@
$\Delta\left(t,\frac{\mathrm{d}^{N}\mathbf{y}}{\mathrm{d}t^{N}},\frac{\mathrm{d}^{N-1}\mathbf{y}}{\mathrm{d}t^{N-1}},\ldots,\dfrac{\mathrm{d}^{2}\mathbf{y}}{\mathrm{d}t^{2}},\dfrac{\mathrm{d}\mathbf{y}}{\mathrm{d}t},\mathbf{y},\boldsymbol{\theta}\right)=\mathbf{0},\\$ (4)
for some power $N\in\mathbb{N}_{+}$ and where $\Delta$ is a vector-valued function of multivariate polynomials. The output system in Eq. (4) is output-equivalent to the original state-output system in Eq. (3) meaning that any outputs $\mathbf{y}$ that are solutions of the latter system are also solutions of the former system and vice versa. The standard differential algebra approach for elucidating structural identifiability analyses the output system, and thus requires that the functions $\mathbf{f}$ and $\mathbf{h}$ defining the state-output system be rational functions of the states. Unlike the standard differential algebra approach, the approach for analysing structural identifiability based on Lie symmetries does not depend on these assumptions. However, since the the aim of this work is to establish the link between these two approaches we restrict ourselves to rational functions $\mathbf{f}$ and $\mathbf{h}$. Provided this problem formulation, we next present the definition of global and local structural identifiability as well as the standard differential algebra approach for conducting a global structural identifiability analysis.
2.1 Global and local structural identifiability
We first present the definition of global structural identifiability [16, 17, 18].
@@ 2.1 Global and local structural identifiability
An individual parameter $\theta_{j}\in\boldsymbol{\theta}$, ${j}\in{1,\ldots,p}$ is globally structurally identifiable if for almost every value $\boldsymbol{\theta}^{\star}$ and for almost all initial conditions the following holds: